Given the DE: y"-(x+1)y'-y=0 use it to answer the following:
a) Find the singular point(s), if any, and if lower bound for the radius of convergence for a power series solution about the ordinary points x=0
b)The recurrence relation Hint: It will be a 3-term recurrence relation
c)Give the first four non-zero terms of each of the two linearly independent power series solutions near the ordinary point x=0
Answer #1

Solution : Siven, y - JOC+ 1)y! -y=0 (1 Now compare with y+ P(x):Y+916) Y=0. → D (OC) = - (C+1) & 26%) = -1 since Play & Q= P2280 - Q4 seº-2020+ [lo+2)(n+1 Jan 72-(n+1) An- lon+2 )An+1]&#07 que compare the coefficients of powers of e both sides,= BOUL+20610 +189 +800 4894 + 2800 6! 61 § 27 = ag+Q6 - 18Q+ +880 + 4894 +2890 - S! - 61 7. = 10821+ 4890 + 4897 +2800 = 1569